Shift report 20011229 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Dec 29 2001 - 15:50:53 EST

  • Next message: Apache: "Shift report 20011230 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Dec 29 15:50:53 2001, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#11 pages 112 - 116
    
    Shift Summary:
    ------------------------------------------------------------
    -- Runs
    
    Run 6193  trig 3 9242  trig 6 35223
    
    8 deg (1/4 field)
      
    Run 6194  trig 3: 35,290 trig 6:95,507
    Run 6195  trig 3: 10,510 trig 6:28,301
    
    
    -- Beam hours so far
    	8-10:10     2 hours
    	12:20-14:20 2 hours
            14:20-15:20 1 hour
    
    	Total Beam this shift 5.25 hours
    
    -- problems
    	T3-A3 (anode) sometimes end up drawing lot of current. (Corona)
    	To fix turn Voltage down to ~ 1100V and turn back up.9:15. Went to physics and picked up the previous log books. The
    	settings in Au run towards the end was actually a bit hard to find
          T5 documented on book10 -page 21 (current values as was yesterday.)
          Later on page 66 (values for T4,T5 says +10V starting run 5544)
     
    	The HV check list from Nov23 says
    	T5: 1700,1685,1595
            (set on 11/6) log book 10- page 107.
    
    10:00 Stop run 6193.
    
    10:05 The t3 anode 3 (v planes) recorded a high current ( 24 micro amps). This is 
    	a clear corona effect. Turn down to 11 and then back to 1190 and current is only
    	1 micro amp.
    	Dana reported the same effect yesterday.
    
    10:10 Seems MCR wants to dump. Turn anodes down.
    
    10:22 Counted beam hours previous shift 0-8 (1h+2.25) 3.25 hours	
    10:25 Access (1h phenix/star) scheduled for 16:00.
    
    10:35 
    	We now have 275K trigger 6 for 8 deg 1/8 field.
    	Deemed enough.
    	Changed FS magnetic field to 1/4B.(D1 - 844)
    
    	Collect ~200 K trigger 6 -> 400K trigger 6 at A pol.
    	Do a zero field run between polarity settings.
    
    10:45   Editted the T3,T4 and T5setting file to reflect changes setting value
            to those used at end of HI run.
    
    11.40	Finished adding a 16 channle discr to the Right Inel counters.
    	Cabled up. Ready for checkout once beam is back.
    
    11:50   Still at injection. The refilling is fairly slow and not great.
    
    12:15	Beam stored and cogged. Pol measurement will start.
    	Inel rate is ~6,000 very good rate - the best so far in this run.
    	The main reason is obviously it is NOT polarized!!!
    
    12:18	Run 6194 begin
    	changed scaledown (inel rate ~6,000/sec).
    
    12:35	The last Left inel tdc's have yet to be cabled up properly.
    	So for run 6194 (L4(1-4) and L3(4) tdc's are not in data stream.
    12:40	All except the L4-4 is in.
    	Supermon still segments violtes (due to inel stuff probably).
    	This is now fixed, but not understood]
    
    14:24   Stop run 6194 for 2 hour limit	
    14:25	Start run 6195
    	Initial inel rate ~4500/sec.i
    15:15 	MCR call to declare pol measurement (it is 0)
    15:16	Stop run 6195
    15:20	Beam dumped for a test ramp, access at 16:30 (Star + Phenix)
    15:30	Pedestal run 6197 (was labbled as data run)
    15:45	TPC pedestal run 6198
    
    	The trigger summary selection still has problems.
    
    	The printing from supermon id delayed. The last ~10 pages only
    	appears next time you print.	
    --other accomplishments
    
    Trigger/scaler setup
    Confirmed setup of Bunch Blue1 + Yellow 21 and correct picking of filled
    overlap bunch for inel overlap.
    
    Updated run check sheet to reflect PP running (renamed too).
    
    Modified defaultmap.txt to get rid(in pp running) of the mapped 0-3, 122-124
    channels for Tofw
    
    
    ****** detailed log for details ********************
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Dec 29 2001 - 15:51:28 EST